I am trying to find a answer to this word problem: Researchers were trying to find how much fish is in the lake so, they caught 200 and tagged them. The next time the went they caught 250 and out of that 16 was tagged. so how many fish is in the lake
1
Expert's answer
2012-09-07T11:45:17-0400
Let N is the fish population.
200/N is the percentage of tagged
ones.
200/N = 16/250, so N = 200 * 250 / 16 = 3125.